OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later. A. 53 B. 63 C. 61 D. 59

OpenStudy (mysesshou):

y = -2x + 65 x=6 y = -2(6) + 65 y= -12 + 65 y=53